3. The summarized financial statements of Baraka enterprises ltd are
As follows
Income statement for the yare ended 30 September
2003 2004
Sh ‘000’ sh’000’
Sales 20,000 28,000
Cost of sales (15,000) (21,000)
Gross profit 5,000 7,000
Administrative expenses 3,800 4,600
Debenture interest - (400)
Net profit 1,200 2,000

Balance sheet as at 30 September
Sh’000’ sh’000’
Non-current assets
Plant, properties and equipment
(NBV) 11,000 14,000
Current assets
Inventories 2,000 3,000
Trade and other receivables 2,500 2,800
Balance at bank - 500
4,500 6,300
Total assets 15,500 20,300
Equity and liabilities
Issued and fully paid
1,000,000 ordinary shares of sh 10 each 10,000 10,000
Revenue reserves 3,000 4,100


Noncurrent liabilities
8% debentures - 5,000
Current liabilities
Trade and other payable 1,500 1,200
Bank overdraft 1,000 -

Stock as at 1 October 2002 was shs 5,000,000

Required:
a) for each year calculate the following
a) Gross profit margin. (2 Marks)
b) Inventory turnover. (2 Marks)
c) Return on equity. (2 Marks)
d) Return on assets. (2 Marks)
e) Acid test ratio. (2 Marks)
f) Current ratio. (2 Marks)
g) Financial leverage. (3 Marks)

b) Comment on the liquid position of the company giving possible
     reasons for the changes.
Answer #1

A)

a)Gross profit margin = GP/Sales

5000/20000=25% , 7000/28000=25%

b) Inventory turnover= COGS/AVG INVENTORY

15000000/(5000000+2000000)= 4.28

21000000/(2000000+3000000)=8.40

c) Return on Equity = Net profit/Equity Capital

1200/10000= 12%, 2000/10000= 20%

d) Return on assets = Net profit/ Avg Total Assets

1200/15500=7.74, 2000/20300= 9.85

e) Acid Test ratio = ( Current Assets- Inventory)/ Current Liabilities

2500/2500= 1, 3300/1200= 2.75

f) Current Ratio = Current Assets/Current Liabilities

4500/2500= 1.8, 6300/1200= 5.25

g) Financial Leverage = EBIT/EBT

1200/1200=1, 4600/4200= 1.095

B)

The liquid position of the company in the year 2003 is within a good line. The acid test ratio shall be minimum 1:1 so that the company can be said to be in a great positive position. Hence the company has a very positive cash test.